A rectangle has a perimeter of 60 inches. The length of the rectangle is 2 inches more than 3 times its width. What are the dimensions of the rectangle?
Answer by Gogonati(855)

We denote x the width, than its length is: 3x+2, Since the perimeter of the rectangle is 60 inches we write the equation:
2x+2(3x+2)=60 (Because the opposite sides are equal)
2x+6x+4=60
8x+4=60
8x=56
x=7
Answer: the width of rectangle is 7 inches and the length is: 3x7+2=23 inches.